# Compile Scene Graphs with Reinforcement Learning

### v1

@@ Abstract
Next token prediction is the fundamental principle for training large language models (LLMs), and reinforcement learning (RL) further enhances their reasoning performance. As an effective way to model language, image, video, and other modalities, the use of LLMs for end-to-end extraction of structured visual representations, such as scene graphs, remains underexplored. It requires the model to accurately produce a set of objects and relationship triplets, rather than generating text token by token. To achieve this, we introduce R1-SGG, a multimodal LLM (M-LLM) initially trained via supervised fine-tuning (SFT) on the scene graph dataset and subsequently refined using reinforcement learning to enhance its ability to generate scene graphs in an end-to-end manner. The SFT follows a conventional prompt-response paradigm, while RL requires the design of effective reward signals. Given the structured nature of scene graphs, we design a graph-centric reward function that integrates node-level rewards, edge-level rewards, and a format consistency reward. Our experiments demonstrate that rule-based RL substantially enhances model performance in the SGG task, achieving a zero failure rate—unlike supervised fine-tuning (SFT), which struggles to generalize effectively. Our code is available at https://github.com/gpt4vision/R1-SGG.
Refer to caption $\max\mathbb{E}_{(I,G)\sim\mathcal{D}}{P(G\mid I)}$ (a) Traditional SGG methods directly maximize the expectation of the likelihood $\mathbb{E}_{(I,G)\sim\mathcal{D}}{P(G\mid I)}$, where image-graph pairs $(I,G)$ are sampled from the dataset $\mathcal{D}$. Refer to caption $\max\mathbb{E}_{(I,G)\sim\mathcal{D}}[$$P(w_{t}\mid w_{0}w_{1}\cdots w_{t-1})]$ (b) M-LLM with SFT is optimized token by token (here, $w_{i}$ refers to a token). Refer to caption $\max\mathbb{E}_{(I,G)\sim\mathcal{D}}[\text{Reward}(V,\hat{V})$$+\ \text{Reward}(E,\hat{E})$+]$\cdots$]+ ⋯ ] (c) M-LLM with RL is optimized using rule-based rewards. Here, $G=(V,E)$ and $\hat{G}=(\hat{V},\hat{E})$ refer to the ground-truth and predicted scene graphs, respectively.
Figure 1: Comparison of traditional Scene Graph Generation (SGG), multimodal LLMs (M-LLMs) with supervised fine-tuning (SFT), and M-LLMs with reinforcement learning (RL) for SGG.
@@ 1 Introduction
Scene graphs, as structured visual representations, have gained increasing attention in many vision applications, such as robot manipulation (Zhu et al., 2021; Zhang et al., 2025), robot navigation (Gu et al., 2024; Miao et al., 2024; Yin et al., 2024), and medical image/video analysis (Lin et al., 2022; Özsoy et al., 2022), etc. To generate scene graphs from an image, traditional Scene Graph Generation (SGG) models (Johnson et al., 2015; Xu et al., 2017; Li et al., 2017; Zellers et al., 2018; Tang et al., 2019; Chen et al., 2019; Li et al., 2022a; Khandelwal and Sigal, 2022; Cong et al., 2023; Zhang et al., 2023; Chen et al., 2024) decouple the task into two subtasks, i.e., object detection and visual relationship recognition, and directly maximize the likelihood of the ground-truth labels given the image (Fig. 1 (a)). Essentially, these models tend to overfit the distribution of annotated datasets; consequently, they struggle to handle long-tail distributions and are prone to generating biased scene graphs (e.g., all predicted relationships are head classes like “on” and “of”).
@@
While traditional SGG models rely on manual annotated datasets and struggle to generalize to new domains, recent advances in large languge models (LLMs) offer a new paragdim. LLM4SGG (Kim et al., 2023) utilizes an LLM to extract relationship triplets from captions using both original and paraphrased text, while GPT4SGG (Chen et al., 2023) employs an LLM to synthesize scene graphs from dense region captions. Additionally, Li (Li et al., 2024) generates scene graphs via image-to-text generation using vision-language models (VLMs). These weakly supervised methods demonstrate potential for generating scene graphs with little or no human annotation but suffer from accuracy issues in the generated results.
@@
Despite these advancements, existing methods typically employ text-only LLMs or rely on intermediate captions as input, which may not fully leverage the rich visual context. In contrast, multimodal large language models (M-LLMs) which integrate both visual and linguistic modalities offer the potential for more direct and holistic scene understanding. By processing visual information alongside natural language prompts, M-LLMs can generate scene graphs in an end-to-end manner. However, in practice, M-LLMs suffer from instruction following (e.g., the output does not contain “objects” or “relationships”), repeated response (e.g., "objects”:[… "id": "desk.9", "bbox": [214, 326, 499, 389], "id": "desk.10", "bbox": [214, 326, 499, 389], "id": "desk.11", "bbox": [214, 326, 499, 389], …]), inaccurate location, etc. These challenges highlight the need for better alignment between visual understanding and structured representation within the M-LLM framework.
@@
To improve instruction-following and structured output generation in M-LLMs, one intuitive solution is to perform Supervised Fine-tuning (SFT) on scene graph datasets (see Fig. 1) (b)). In the context of SGG, SFT aligns the model’s outputs with expected formats (e.g., structured lists of objects and relationships) by training it on high-quality scene graph annotations. This process encourages the model not only to recognize entities and relations from the image but also to organize them into a coherent and valid graph structure. Nevertheless, SFT alone may still be insufficient as all output tokens are weighted equally in the loss. For example, the experimental results on the VG150 dataset Xu et al. (2017) reveal that even with SFT, M-LLM still has a high failure rate to generate a valid and high-quality scene graph. The drawback of SFT in SGG lies in the lack of effective signals to correct the output (e.g., the model cannot directly utilize the Intersection over Union (IoU) between the predicted box and the ground truth to refine its output ).
@@
To advance M-LLMs for effective Scene Graph Generation (SGG), we propose R1-SGG, a novel framework leveraging visual instruction tuning enhanced by reinforcement learning (RL). The visual instruction tuning stage follows a conventional supervised fine-tuning (SFT) paradigm, i.e., fine-tuning the model using prompt-response pairs with a cross-entropy loss. For the RL stage, we adopt GRPO, an online policy optimization algorithm introduced in DeepSeekMath Shao et al. (2024).
To effectively apply GRPO to the SGG task, we introduce several rule-based rewards, reflecting key characteristics of scene graphs. Generally, given an image and an associated textual prompt, the multimodal LLM generates a set of predicted objects and relationship triplets. To establish a precise node-level correspondence between the predicted set and the ground-truth annotations, we formulate this alignment problem as a bipartite matching problem, efficiently resolved using the Hungarian algorithm (Kuhn, 1955). Upon solving the bipartite matching, we acquire an optimal mapping between predicted nodes and ground-truth nodes. The rewards are then systematically computed as follows: 1) Node-level reward: Calculated as the embedding similarity between the predicted object and the corresponding ground-truth object, combined with the Intersection-over-Union (IoU) of their bounding boxes. 2) Edge-level reward: Measured by the embedding similarity of the predicted predicate and the ground-truth predicate associated with matched object pairs, reflecting the accuracy of relationship predictions. This graph-centric reward design along with a format reward (e.g., the output should contain “objects” and “relationships”) effectively guides the policy optimization process to refine the model for generating accurate, contextually consistent, and diverse scene graphs.
@@ 2 Related Work
Scene Graph Generation (SGG) is a foundational task in structured visual understanding, where the goal is to represent an image as a graph of objects and their pairwise relationships. Traditional approaches like Xu et al. (2017); Li et al. (2017); Zellers et al. (2018); Tang et al. (2019); Chen et al. (2019); Li et al. (2022a); Khandelwal and Sigal (2022); Cong et al. (2023) decouple the task into object detection and relationship classification stages, and are typically trained via supervised learning on datasets such as Visual Genome (VG150) (Xu et al., 2017). While effective, these models are limited by their reliance on annotated data and exhibit strong bias toward head predicates such as “on” or “of”, struggling on long-tail classes.
To overcome the closed-set limitation, recent work has explored open-vocabulary SGG. For example, OvSGTR (Chen et al., 2024) extends scene graph prediction to a fully open-vocabulary setting by leveraging visual-concept alignment. In parallel, weakly supervised methods have been developed to reduce the annotation burden. These approaches, such as those proposed by Zhong et al. (2021); Li et al. (2022b); Zhang et al. (2023); Chen et al. (2024), use image-caption pairs as supervision to distill relational knowledge, enabling generalization to unseen concepts.
@@
With the rise of LLMs, several studies have attempted to synthesize scene graphs from natural language. LLM4SGG (Kim et al., 2023) extracts relational triplets from both original and paraphrased captions using text-only LLMs. GPT4SGG (Chen et al., 2023) goes a step further by using GPT-4 to generate scene graphs from dense region captions, improving contextual consistency and coverage. Meanwhile, Li et al. (2024) leverage vision-language models (VLMs) to produce scene graphs through image-to-text generation pipelines.
@@
Reinforcement learning (RL) has been increasingly adopted to enhance the reasoning capabilities of large models. Algorithms like Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) guide models using reward signals instead of relying solely on maximum likelihood estimation. In the context of large language models, DeepSeek-R1 (Guo et al., 2025) demonstrates that RL can significantly improve structured reasoning and planning.
@@ 3.1 Preliminary
Group Relative Policy Optimization (GRPO) is a online reinforcement learning algorithm introduced by DeepSeekMath (Shao et al., 2024). Unlike traditional methods such as PPO (Schulman et al., 2017), which require an explicit critic network, GRPO instead compares groups of candiates to update the policy $\pi_{\theta}$. Specifically, for each input query $q$, a set of candidate outputs $\{o_{i}\}_{i=1}^{G}$ is drawn from the previous policy $\pi^{\text{old}}(O|q)$, and the advantage of each candidate is computed relative to the group’s average reward:
@@
Here, $\epsilon$ and $\beta$ are hyper-parameters. The first term uses a clipped probability ratio (as in PPO) to control the update magnitude, while the KL divergence regularizer $D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})$ constrains the new policy $\pi_{\theta}$ to not deviate too much from a reference policy $\pi_{\text{ref}}$. This formulation, which combines a group-relative advantage, a clipping mechanism, and a KL divergence regularizer, stabilizes policy updates and improves training efficiency, demonstrating remarkable potential for enhancing the reasoning performance of LLMs such as DeepSeek R1 (Guo et al., 2025).
@@ 3.2 Overview of R1-SGG
Graph Matching.
Given an image and a text prompt, the M-LLM outputs a directed graph ${\mathcal{G}}_{\text{pred}}$, which needs to be matched to the ground-truth graph $\mathcal{G}$. To achieve this, we perform bipartite matching, as used in DETR (Carion et al., 2020), to associate the prediction set $\{v_{i}\}_{i=1}^{M}$ with the ground-truth set $\{\tilde{v}_{i}\}_{i=1}^{N}$. The cost objective of the matching is formulated as:
@@
3.3 Rewards Definition
Format Reward.
Following DeepSeek R1 (Guo et al., 2025), we employ a format reward to ensure that the model’s response adheres to the expected structure, specifically <think>$\cdots$</think><answer>$\cdots$</answer>. A reward of 1 is assigned if the response follows this format and the segment enclosed by <answer>$\cdots$</answer> contains both the keywords "object" and "relationships"; otherwise, the reward is 0. However, since small models such as Qwen/Qwen2-VL-2B-Instruct have limited ability to follow the system prompt, we relax this constraint by awarding a reward of 0.5 for responses that are formatted as valid JSON dictionaries, even if they do not strictly meet the original criteria.
Node-level Reward.
@@ 4.1 Dataset and Experiment Setup
Dataset.
Table 1: Prompting an M-LLM to generate scene graphs without providing predefined object classes or predicate types. The version with predefined classes and predicates is available in our code repository.
messages = [{ "role": "system", "content": " {system_prompt}" }, { "role": "user", "content": f"""Generate a structured scene graph for an image of size using the following format: “‘json { "objects": [ {"id": "object_name.number", "bbox": [x1, y1, x2, y2]}, … ], "relationships": [ {"subject": "object_name.number", "predicate": "relationship_type", "object": "object_name.number"}, … ] }“‘. ### **Guidelines:** - **Objects:** - Assign a unique ID for each object using the format "object_name.number" (e.g., "person.1", "bike.2"). - Provide its bounding box ‘[x1, y1, x2, y2]’ in integer pixel format. - Include all visible objects, even if they have no relationships. - **Relationships:** - Represent interactions accurately using "subject", "predicate", and "object". - Omit relationships for orphan objects. ### **Example Output:** “‘json { "objects": [ {"id": "person.1", "bbox": [120, 200, 350, 700]}, {"id": "bike.2", "bbox": [100, 600, 400, 800]}, {"id": "helmet.3", "bbox": [150, 150, 280, 240]}, {"id": "tree.4", "bbox": [500, 100, 750, 700]} ], "relationships": [ {"subject": "person.1", "predicate": "riding", "object": "bike.2"}, {"subject": "person.1", "predicate": "wearing", "object": "helmet.3"} ] } “‘ Now, generate the complete scene graph for the provided image: """ } ]
We leverage the widely-used scene graph dataset, VG150 Xu et al. (2017), which consists of $150$ object categories and $50$ relationship categories. The training set contains $56,224$ image-graph pairs, while the validation set includes $5,000$ pairs. To prompt the M-LLM, we transform the image-graph pair using the template described in Table 1.
Evaluation. Following standard evaluation pipeline in SGG, we adopt the SGDET protocol to measure the model’s ability to generate scene graphs.
Implementation Details. Our code is based on the trl library von Werra et al. (2020) and utilizes vLLM Kwon et al. (2023) to speed up sampling during reinforcement learning. For supervised fine-tuning (SFT), the model is trained for 3 epochs using 4x NVIDIA A100 (80GB) GPUs. For reinforcement learning (RL), the 2B model is trained for 1 epoch using 16x NVIDIA A100 (80GB) GPUs and the 7B model is trained for 1 epoch using 16x NVIDIA GH200 (120GB) GPUs.

### v3

@@ Abstract
Next token prediction is the fundamental principle for training large language models (LLMs), and reinforcement learning (RL) further enhances their reasoning performance. As an effective way to model language, image, video, and other modalities, the use of LLMs for end-to-end extraction of structured visual representations, such as scene graphs, remains underexplored. It requires the model to accurately produce a set of objects and relationship triplets, rather than generating text token by token. To achieve this, we introduce R1-SGG, a multimodal LLM (M-LLM) initially trained via supervised fine-tuning (SFT) on the scene graph dataset and subsequently refined using reinforcement learning to enhance its ability to generate scene graphs in an end-to-end manner. The SFT follows a conventional prompt-response paradigm, while RL requires the design of effective reward signals. We design a set of graph centric rewards, including three recall based variants—Hard Recall, Hard Recall+Relax, and Soft Recall—which evaluate semantic and spatial alignment between predictions and ground truth at the object and relation levels. A format consistency reward further ensures that outputs follow the expected structural schema. Extensive experiments on the VG150 and PSG benchmarks show that R1-SGG substantially reduces failure rates and achieves strong performance in Recall and mean Recall, surpassing traditional SGG models and existing multimodal language models. Our code is available at https://github.com/gpt4vision/R1-SGG.
@@ 1 Introduction
Scene graphs, as structured visual representations, have gained increasing attention in many vision applications, such as robot manipulation [44, 41], robot navigation [7, 23, 37], and medical image / video analysis [20, 24], etc. To generate scene graphs from an image, traditional Scene Graph Generation (SGG) models [10, 34, 14, 38, 29, 2, 15, 11, 5, 40, 4] decouple the task into two subtasks, i.e., object detection and visual relationship recognition, and directly maximize the likelihood of the ground-truth labels given the image (Fig. 1 (a)). Essentially, these models tend to overfit the distribution of annotated datasets; consequently, they struggle to handle long-tail distributions and are prone to generating biased scene graphs (e.g., all predicted relationships are head classes like “on” and “of”).
@@
While traditional SGG models rely on manual annotated datasets and struggle to generalize to new domains, recent advances in large languge models (LLMs) offer a new paragdim. LLM4SGG [12] utilizes an LLM to extract relationship triplets from captions using both original and paraphrased text, while GPT4SGG [3] employs an LLM to synthesize scene graphs from dense region captions. Additionally, Li [17] generates scene graphs via image-to-text generation using vision-language models (VLMs). These weakly supervised methods demonstrate potential for generating scene graphs with little or no human annotation but suffer from accuracy issues in the generated results.
Refer to caption $\max\mathbb{E}_{(I,G)\sim\mathcal{D}}[$$P(w_{t}\mid w_{0}w_{1}\cdots w_{t-1})]$ (a) M-LLM with SFT is optimized token by token (here, $w_{i}$ refers to a token). Refer to caption $\max\mathbb{E}_{(I,G)\sim\mathcal{D}}[\text{Reward}(V,\hat{V})$$+\ \text{Reward}(E,\hat{E})$+]$\cdots$]+ ⋯ ] (b) M-LLM with RL is optimized using rule-based rewards. Here, $G=(V,E)$ and $\hat{G}=(\hat{V},\hat{E})$ refer to the ground-truth and predicted scene graphs, respectively.
Figure 1: Comparison of multimodal LLMs (M-LLMs) fine-tuned via Supervised Fine-tuning (SFT) and Reinforcement Learning (RL) for Scene Graph Generation (SGG).
@@
Despite these advancements, existing methods typically employ text-only LLMs or rely on intermediate captions as input, which do not fully leverage the rich visual context. In contrast, multimodal large language models (M-LLMs) which integrate both visual and linguistic modalities offer the potential for more direct and holistic scene understanding. By processing visual information alongside natural language prompts, M-LLMs can generate scene graphs in an end-to-end manner. However, in practice, M-LLMs suffer from instruction following (e.g., the output does not contain “objects” or “relationships”), repeated response (e.g., "objects”:[… "id": "desk.9", "bbox": [214, 326, 499, 389], "id": "desk.10", "bbox": [214, 326, 499, 389], "id": "desk.11", "bbox": [214, 326, 499, 389], …]), inaccurate location, etc. These challenges highlight the need for better alignment between visual understanding and structured representation within the M-LLM framework.
@@
To improve instruction-following and structured output generation in M-LLMs, one intuitive solution is to perform Supervised Fine-tuning (SFT) on scene graph datasets (see Fig. 1) (b)). In the context of SGG, SFT aligns the model’s outputs with expected formats (e.g., structured lists of objects and relationships) by training it on high-quality scene graph annotations. This process encourages the model not only to recognize entities and relations from the image but also to organize them into a coherent and valid graph structure. Nevertheless, SFT alone still be insufficient as all output tokens are weighted equally in the loss. For example, the experimental results on the VG150 dataset [34] reveal that even with SFT, M-LLM still has a high failure rate to generate a valid and high-quality scene graph. The drawback of SFT in SGG lies in the lack of effective signals to correct the output (e.g., the model cannot directly utilize the Intersection over Union (IoU) between the predicted box and the ground truth to refine its output ).
@@
To advance M-LLMs for effective Scene Graph Generation (SGG), we propose R1-SGG, a novel framework leveraging visual instruction tuning enhanced by reinforcement learning (RL). The visual instruction tuning stage follows a conventional supervised fine-tuning (SFT) paradigm, i.e., fine-tuning the model using prompt-response pairs with a cross-entropy loss. For the RL stage, we adopt GRPO, an online policy optimization algorithm introduced in DeepSeekMath [28].
To enable effective reinforcement learning for Scene Graph Generation (SGG), we introduce a set of rule-based, graph-centric rewards that reflect the structural characteristics of scene graphs. Given an image and a prompt, a multimodal large language model (M-LLM) generates a set of objects and relational triplets. To evaluate and optimize these predictions, we formulate reward functions aligned with standard SGDET metrics [34] and structured reasoning objectives. Specifically, we define three reward variants: Hard Recall, which counts a triplet as correct only if the subject, predicate, and object labels exactly match the ground truth and both bounding boxes achieve IoU > 0.5; Hard Recall+Relax, which relaxes the exact match constraint by incorporating embedding similarity between predicted and ground-truth labels; and Soft Recall, which further densifies reward signals via bipartite matching, combining object label similarity, IoU, and bounding box distance into a unified cost function. These scene graph rewards are computed over matched object and edge pairs, and are complemented by a format reward that enforces structural adherence in output formatting. This reward design enables stable and fine-grained policy optimization using GRPO, guiding the M-LLM toward generating accurate, complete, and structurally valid scene graphs.
@@ 2 Related Work
Scene Graph Generation (SGG) is a foundational task in structured visual understanding, where the goal is to represent an image as a graph of objects and their pairwise relationships. Traditional approaches like [34, 14, 38, 29, 2, 15, 11, 5] decouple the task into object detection and relationship classification stages, and are typically trained via supervised learning on datasets such as Visual Genome (VG150) [34]. While effective, these models are limited by their reliance on annotated data and exhibit strong bias toward head predicates such as “on” or “of”, struggling on long-tail classes.
To overcome the closed-set limitation, recent work has explored open-vocabulary SGG. For example, OvSGTR [4] extends scene graph prediction to a fully open-vocabulary setting by leveraging visual-concept alignment. In parallel, weakly supervised methods have been developed to reduce the annotation burden. These approaches, such as those proposed by [43, 16, 40, 4], use image-caption pairs as supervision to distill relational knowledge, enabling generalization to unseen concepts.
@@
With the rise of LLMs, several studies have attempted to synthesize scene graphs from natural language. LLM4SGG [12] extracts relational triplets from both original and paraphrased captions using text-only LLMs. GPT4SGG [3] goes a step further by using GPT-4 to generate scene graphs from dense region captions, improving contextual consistency and coverage. Meanwhile, [17] leverage vision-language models (VLMs) to produce scene graphs through image-to-text generation pipelines.
@@
Reinforcement learning (RL) has been increasingly adopted to enhance the reasoning capabilities of large models. Algorithms like Proximal Policy Optimization (PPO) [27] and Group Relative Policy Optimization (GRPO) [28] guide models using reward signals instead of relying solely on maximum likelihood estimation. In the context of large language models, DeepSeek-R1 [8] demonstrates that RL can significantly improve structured reasoning and planning.
@@ 3.1 Preliminary
Group Relative Policy Optimization (GRPO) is a online reinforcement learning algorithm introduced by DeepSeekMath [28]. Unlike traditional methods such as PPO [27], which require an explicit critic network, GRPO instead compares groups of candiates to update the policy $\pi_{\theta}$. Specifically, for each input query $q$, a set of candidate outputs $\{o_{i}\}_{i=1}^{G}$ is drawn from the previous policy $\pi^{\text{old}}(O|q)$, and the advantage of each candidate is computed relative to the group’s average reward:
@@
Here, $\epsilon$ and $\beta$ are hyper-parameters. The first term uses a clipped probability ratio (as in PPO) to control the update magnitude, while the KL divergence regularizer $D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})$ constrains the new policy $\pi_{\theta}$ to not deviate too much from a reference policy $\pi_{\text{ref}}$. This formulation, which combines a group-relative advantage, a clipping mechanism, and a KL divergence regularizer, stabilizes policy updates and improves training efficiency, demonstrating remarkable potential for enhancing the reasoning performance of LLMs such as DeepSeek R1 [8].
@@ 3.2 Overview of R1-SGG
R1-SGG is a reinforcement learning framework that enhances scene graph generation (SGG) in multimodal large language models (M-LLMs). It builds on a supervised fine-tuning (SFT) stage using prompt-response pairs, followed by reinforcement learning (RL) with structured, graph-centric rewards.
Given an input image and prompt, the M-LLM generates a scene graph $\mathcal{G}_{\text{pred}}=(\mathcal{V}_{\text{pred}},\mathcal{E}_{\text{pred}})$, comprising objects (nodes) and their relationships (edges). We primarily optimize using Hard Recall, which aligns with SGDET metrics by rewarding exact triplet matches. To study the sparsity and design of the rewards, we also evaluated relaxed alternatives based on bipartite matching between $\mathcal{G}_{\text{pred}}$ and the ground truth graph $\mathcal{G}_{\text{gt}}$, allowing fine-grained node and edge rewards. Our RL pipeline employs Group Relative Policy Optimization (GRPO) [28], which compares sampled outputs and promotes higher-reward candidates. By integrating SFT, GRPO, and graph-aware rewards, R1-SGG enables M-LLMs to generate accurate, diverse, and structurally valid scene graphs.
3.3 Rewards Definition
3.3.1 Format Reward
Following DeepSeek R1 [8], we employ a format reward to ensure that the model’s response adheres to the expected structure, specifically <think>$\cdots$</think><answer>$\cdots$</answer>. A reward of 1 is assigned if the response follows this format and the segment enclosed by <answer>$\cdots$</answer> contains both the keywords "object" and "relationships"; otherwise, the reward is 0.
3.3.2 Scene Graph Rewards
Standard evaluation protocols for Scene Graph Generation (SGG), such as SGDET [34], formulate the task as a recall-oriented problem, emphasizing the model’s ability to retrieve correct relationship triplets from an image. To investigate the impact of different reward formulations, we introduce three variants: Hard Recall, Hard Recall+Relax, and Soft Recall.
Hard Recall.
[…]
This reward is aligned with standard metrics but suffers from sparsity due to its strict criteria.
Hard Recall + Relax.
We relax the triplet accuracy requirement by computing cosine similarity between the entity embeddings of predicted and ground-truth triplets. This softens the discrete matching constraint to provide more gradient signal.
Soft Recall.
We further propose a dense matching reward by formulating it as a bipartite matching problem, similar to DETR [1], where predicted nodes $\{v_{i}=(c_{i},b_{i})\}_{i=1}^{M}$ (each node $v_{i}$ is comprising an object class $c_{i}$ and a bounding box $b_{i}$) are matched to ground-truth nodes $\{\tilde{v}_{j}=(\tilde{c}_{j},\tilde{b}_{j})\}_{j=1}^{N}$ with the following cost:
@@ 4.1 Dataset and Experiment Setup
4.1.1 Dataset.
VG150. The widely-used scene graph dataset VG150 [34] consists of 150 object categories and 50 relation categories. Following prior works [40, 4], the training set used in this work contains 56,224 image-graph pairs, while the validation set includes 5,000 pairs. To prompt the M-LLM, we transform each image-graph pair using the template described in Table 6.
PSG. The Panoptic Scene Graph (PSG) dataset [36] is built on the COCO dataset [18], consisting of 80 thing object categories, 53 stuff object categories, and 56 relation categories. It contains 46,563 image-graph pairs for training and 2,186 pairs for testing.
4.1.2 Evaluation
Following the standard evaluation pipeline in SGG, we adopt the SGDET protocol [34, 30] to measure the model’s ability to generate scene graphs. SGDET requires the model to generate scene graphs directly from the image without any predefined object boxes. Performance is evaluated using Recall and mean Recall (mRecall). Recall is computed for each image-graph pair, where a predicted triplet is considered correct if both the subject and object bounding boxes have an Intersection over Union (IoU) of at least 0.5 with the corresponding ground-truth boxes, and the subject category, object category, and relationship label all match the ground truth. Mean Recall (mRecall) is obtained by averaging the Recall across all relation categories. We additionally report AP@50 to assess object detection performance and Failure Rate to evaluate format consistency.
4.1.3 Implementation Details
Our code is based on the trl library [31] and utilizes vLLM [13] to speed up sampling during reinforcement learning. For SFT, the model is trained for 3 epochs with a batch size of 128 on 4 NVIDIA A100 (80GB) GPUs, using the AdamW optimizer [22] with a maximum learning rate of 1e-5. For RL, the model is trained for 1 epoch with a batch size of 32 and 8 generations per sample on 16 NVIDIA GH200 (120GB) GPUs, also using AdamW with a maximum learning rate of 6e-7.